# Zhaojian Yu, Xin Zhang, Ning Shang, Yangyu Huang, Can Xu, Yishujie Zhao, Wenxiang Hu, Qiufeng Yin Microsoft

### 26 Dec 2023 v2

@@ Abstract
Recent work demonstrates that, after being fine-tuned on a high-quality instruction dataset, the resulting model can obtain impressive capabilities to address a wide range of tasks. However, existing methods for instruction data generation often produce duplicate data and are not controllable enough on data quality. In this paper, we extend the generalization of instruction tuning by classifying the instruction data to 4 code-related tasks and propose a LLM-based Generator-Discriminator data process framework to generate diverse, high-quality instruction data from open source code. Hence, we introduce CodeOcean, a dataset comprising 20,000 instruction instances across 4 universal code-related tasks,which is aimed at augmenting the effectiveness of instruction tuning and improving the generalization ability of fine-tuned model. Subsequently, we present WaveCoder, a fine-tuned Code LLM with Widespread And Versatile Enhanced instruction tuning. This model is specifically designed for enhancing instruction tuning of Code Language Models (LLMs). Our experiments demonstrate that Wavecoder models outperform other open-source models in terms of generalization ability across different code-related tasks at the same level of fine-tuning scale. Moreover, Wavecoder exhibits high efficiency in previous code generation tasks. This paper thus offers a significant contribution to the field of instruction data generation and fine-tuning models, providing new insights and tools for enhancing performance in code-related tasks.
@@ 1 Introduction
Figure 2: The overview of the pipeline.
@@
Additionally, instruction tuning has been shown to perform at a satisfactory level across various tasks, as demonstrated in research based on instruction-tuning like FLAN [8], ExT5 [9], and FLANT5 [10]. These studies incorporated thousands of tasks into their training pipelines to improve the generalization ability of pre-trained model on downstream tasks. InstructGPT [11] aligns user input effectively by integrating high-quality instruction data written by human labelers, which inspires further exploration into the utility of instruction tuning. Stanford Alpaca [12] utilized ChatGPT to generate instruction data via a self-instruct method, which in turn was used in the instruction tuning process. Further improvements to the effectiveness of pre-training models have been explored in WizardLM [13] and WizardCoder [14], where the evol-instruct method was applied. These works collectively exemplify the potential of instruction tuning in enhancing the performance of large language models.
@@
To validate our approach, taking StarCoder [6], CodeLLaMa [7], and DeepseekCoder [22] as our base models , we introduce WaveCoder models, fine-tuned with our data generation strategy and evaluate it on HumanEval [4], MBPP [23], HumanEvalPack [24] benchmarks, experimental results show that our WaveCoder exhibits exceptional performance based on a small-scale instruction tuning.
@@ 2 CodeOcean: Four-task Code-related Instruction Data
Figure 3: The overview of the LLM-based Generator-Discriminator Framework.
@@ 2.1 Tasks Setting
Given the breadth of code-related tasks, we have selected four of the most universally representative and common tasks from the three generative tasks (code-to-text, text-to-code, and code-to-code) for further exploration. These include Code Summarization, Code Generation, Code Translation, and Code Repair. Detailed descriptions of these tasks can be found below.
@@ 2.2 Generation of Training Data
In past research work[25, 26], many researchers have discovered that data quality and diversity play an important role in the instruction tuning process rather than data amount. The improvement of data quality and diversity are directly related to the performance of the fine-tuned model. Therefore, we introduce the details of our training data generation process in this section.
2.2.1 Selection of Raw Code
CodeSearchNet [^2] https://huggingface.co/datasets/code_search_net is a dataset containing 2 million (comment, code) pairs from open-source libraries hosted on GitHub. It includes code and documentation for 6 programming languages. We have chosen CodeSearchNet as our foundation dataset and applied the coreset-based selection method KCenterGreedy [27, 28] to maximize the diversity of the raw code.
In order to select high-quality code for instruction-tuning, we make some rules to filter the foundation dataset preliminarily. The main rules are as follows:
@@ 2.2.1 Selection of Raw Code
After filtered by the above rules, the size of the raw code dataset dropped to 1.2 million, which is still a huge scale and too expensive to fine-tune. To this end, we employed KCenterGreedy algorithm based on code embeddings to reduce the amount of training data while maximizing data diversity of raw code as much as possible.
@@
Table 2: The proportion of generated data and the prompts used.
@@ 2.2.2 LLM-based Generator-Discriminator Framework
After the process of collecting raw code, it is significant to generate instruction data and reformat them to training data. To this end, shown in Figure 2, we propose a novel LLM-based Generator-Discriminator Framework that can leverage an extensive amount of unsupervised open source code to generate supervised instruction data. By incorporating the open source code, the diversity of the generated data no longer relies solely on capability of the teacher LLM itself.
In the generation phase, we utilize GPT-4 to generate a task definition in the context of a specific scenario. We follow this with the definition of generation requirements for the task at hand, integrating both the task definition and all associated requirements into the generation prompt. By using the raw code as an input and selecting good and bad case examples, we leverage the few-shot abilities of GPT-3.5 to generate the necessary knowledge and information for instruction tuning.
For the discriminator phase, we establish a series of criteria related to instruction data and employ GPT-4 to assess the quality of instruction instances. Each instruction instance is classified as either a good or bad case, and this information is subsequently reused in the following generation as examples. An example of our Generator and Discriminator framework is presented in Figure 4. This framework provides a comprehensive approach to generating and evaluating instruction data, ensuring a high-quality and diverse training dataset.
Figure 4: An example of LLM-based Generator and Discriminator framework. The generator produce the instruction data based on the input (a). Subsequently, the discriminator accept the output and generated analysis for it. The output (b) includes four keys, we take the information as input and solution as output in our instruction tuning. The analysis (c) consists of the detailed reason for each rule and an overall answer to check if the sample meet all the requirements.
3 Training
@@
To obtain WaveCoder models, We choose StarCoder-15B, CodeLLaMa (7B and 13B), DeepseekCoder-6.7B as the base model and fine-tune all the base model for 3 epochs using NVIDIA A100-80GB GPU. For StarCoder-15B and CodeLLaMa-13B, we set the global batch size to 128 using Tensor Parallel and set the initial learning rate at 2e-5. For DeepseekCoder-6.7B and CodeLLaMa-7B, we set the global batch size to 256 using the Fully Sharded Data Parallel (FSDP) module from Pytorch and set the initial learning rate at 5e-5. The sequence length of our each instruction sample is less than 1024. The fine-tuning prompt is as follows:
@@
4 Experiments
4.1 Experimental Setup
4.1.1 Benchmarks and Baselines
MBPP  [^5] https://huggingface.co/datasets/mbpp consists of around 1,000 crowd-sourced Python programming problems, designed to be solvable by entry level programmers, covering programming fundamentals, standard library functionality, and so on. In this paper, we choose the 500 preoblems test dataset to evaluate both few-shot inference of fine-tuned models. For whose MBPP(500) result is not reported or not used, we reproduced for them using bigcode-evaluation-harness  [^6] https://github.com/bigcode-project/bigcode-evaluation-harness.
Proprietary Models. We present the self-reported results from an array of SoTA LLMs, including Gemini, ChatGPT (gpt-3.5-turbo), GPT-4. If not reported, we use the results from Octopack or evaluate by ourselves.
4.1.2 metrics
$pass@k=E\left[1-\frac{{n\choose k-c}}{{n\choose k}}\right]$ (1)
4.2 Experimental Results
4.2.1 Evaluation on Code Generation Task
WaveCoder models substantially outperform the instructed models training with fewer than 20K instruction tuning data (InsT Data). Following the fine-tuning process, the performance of our models exhibit substantial improvement when compared to both the foundation model and a selection of open-source models, but it still lags behind proprietary models the instructed models training with more than 70K training data.
Refined and diverse instruction data can significantly improve the efficiency of instruction tuning. Although Wavecoder falls behind WizardCoder and Magicoder on code generation task, there is a considerable difference in data size among them. Nevertheless, the performance lag of Wavecoder is relatively minor.
@@ 4.2.1 Evaluation on Code Generation Task
DeepseekCoder 6.7B - 49.4 60.6
WaveCoder-DS-6.7B [Uncaptioned image] 6.7B DeepseekCoder 20K 62.8 (+13.4) 62.4 (+1.8)
@@
WaveCoder-DS-6.7B [Uncaptioned image] 54.9 51.2 57.3 49.3 45.1 32.9 48.4
@@ 4.2.2 Evaluation on Other Code-related Task
We score WaveCoder with state-of-the-art Code LLMs on HumanEvalPack in Table 4.We use the result reported in Octopack [21] when avaliable. Table 4 presents the results of WaveCoder on code repair task and Table 5 presents the results of WaveCoder on Code summarization task, highlighting the the following salient observations:
@@
WaveCoder-DS-6.7B [Uncaptioned image] 47.5 45.7 48.2 33.5 41.4 27.4 40.6
@@ 4.3 Comparison with CodeAlpaca
  • After being fine-tuned with 1K, 5K and 20K of instructional data respectively, the performance of base model improves significantly on HumanEval shown in Figure 1.Taking Starcoder as the base model, CodeOcean surpasses the CodeAlpaca (44.9% vs 41.7%, 45.7% vs 48.1% and 47.0% vs 50.5%) shown in Figure 1 (a), which emphasizes the effectiveness of our LLM-based Generator-Discriminator Framework on refining instruction data. As shown in Figure 1 (b), The results of different base models on CodeOcean surpasses the the results on CodeAlpaca, which emphasizes the effectiveness of CodeOcean dataset in enhancing the instruction-following ability of the base model.
  • According to Table 4 and Table 5, All Wavecoder models significantly outperform the model fine-tuned with CodeAlpaca. Remarkably, The pass@1 score of Wavecoder-CL-13B outperforms CodeLLaMa-CodeAlpaca-13B achieving 10.6% absolute improvements on HumanEvalExplain. This emphasizes the effectiveness of defining and classifying code-related tasks on enhancing the generalization ability of Code LLMs.
@@ 5 Related Work
5.1 Code Large Language Models.
Numerous researchers have proposed their Code Language Models (LLMs), including CodeGen [32], CodeT5[33], StarCoder [6], CodeLLaMa [7], and Deepseek-Coder [22] within coding domain. By leveraging a comprehensive pre-training process with a vast amount of code knowledge from scratch, these Code LLMs often exceed the performance of super LLMs in many scenarios. In a bid to generate output more efficiently and assist in resolving programming issues, instruction-based models like InstructCodeT5+ [34] and WizardCoder [14] have been introduced.
5.2 Instruction Tuning.
In order to unleash the potential of the foundation LLMs, instruction tuning is proposed to fine-tune the LLM with multi-task datasets which is presented in task descriptions by natural language. For example, OpenAI fine-tuned its GPT-3 [3] model with a diverse instruction dataset written by human annotators, the instruction-tuned model can effectively align the intention of human users. This work is the indispensable part of the InstructGPT [17], which is the foundation of the well-known ChatGPT. [10] proposes Flan-PaLM 540B [10] which is instruction-tuned on 1.8K tasks and finds that the diversity of instruction dataset can significantly improve the performance of the instruction-tuned language model. Alpaca [12] adopts the self-instruct method, which leverages the few shot ability of ChatGPT to generate data for training. For Code LLMs, InstructCodeT5+ [34], WizardCoder [14], Pangu-coder2 [29], are recent recent models trained with coding instructions. However, all the instruction data they used is from CodeAlpaca, which is not refined enough.

### 11 Jan 2024 v3

@@ Abstract
Recent work demonstrates that, after being fine-tuned on a high-quality instruction dataset, the resulting model can obtain impressive capabilities to address a wide range of tasks. However, existing methods for instruction data generation often produce duplicate data and are not controllable enough on data quality. In this paper, we extend the generalization of instruction tuning by classifying the instruction data to 4 code-related tasks which is specifically designed for enhancing instruction tuning of Code Large Language Models (LLMs) and propose a LLM-based Generator-Discriminator data process framework to generate diverse, high-quality instruction data from open source code. Hence, we introduce CodeOcean, a dataset comprising 20,000 instruction instances across 4 universal code-related tasks, which is aimed at augmenting the effectiveness of instruction tuning and improving the generalization ability of fine-tuned model. Subsequently, we present WaveCoder, a fine-tuned Code LLM with Widespread And Versatile Enhanced instruction tuning. Our experiments demonstrate that WaveCoder outperforms other open-source models in terms of generalization ability across different code-related tasks at the same level of fine-tuning scale. Moreover, WaveCoder exhibits high efficiency in previous code generation tasks. This paper thus offers a significant contribution to the field of instruction data generation and fine-tuning models, providing new insights and tools for enhancing performance in code-related tasks.
@@ 1 Introduction
Figure 2: The overview of the pipeline. Part A indicate the collection of raw code. Part B and C indicates the LLM-based Generator and LLM-based Disciminator respectively.
@@
Additionally, instruction tuning has been shown to perform at a satisfactory level across various tasks, as demonstrated in research based on instruction-tuning like FLAN [8], ExT5 [9], and FLANT5 [10]. These studies incorporated thousands of tasks into their training pipelines to improve the generalization ability of pre-trained model on downstream tasks. InstructGPT [11] aligns user input effectively by integrating high-quality instruction data written by human labelers, which inspires further exploration into the utility of instruction tuning. Stanford Alpaca [12] utilized ChatGPT to generate instruction data via a self-instruct method, which in turn was used in the instruction tuning process. Further improvements to the effectiveness of instruction tuning have been explored in WizardLM [13] and WizardCoder [14], where the evol-instruct method was applied. These works collectively exemplify the potential of instruction tuning in enhancing the performance of large language models.
@@
To validate our approach, taking StarCoder [6], CodeLLaMa [7], and DeepseekCoder [22] as our base models, we introduce WaveCoder models, fine-tuned with our data generation strategy and evaluate them on HumanEval [4], MBPP [23], HumanEvalPack [24] benchmarks, experimental results show that our WaveCoder exhibits exceptional performance based on a small-scale instruction tuning.
@@ 2.1 Tasks Setting
Given the breadth of code-related tasks from CodeXGLUE [25], we have selected four of the most universally representative and common tasks from the three generative tasks (code-to-text, text-to-code, and code-to-code) for further exploration. These include Code Summarization, Code Generation, Code Translation, and Code Repair. Detailed descriptions of these tasks can be found below.
@@ 2.2 Generation of Training Data
In past research work [26, 27], many researchers have discovered that data quality and diversity play an important role in the instruction tuning process rather than data amount. The improvement of data quality and diversity are directly related to the performance of the fine-tuned model. Therefore, to ensure the data quality and diversity of instruction instance, we propose: 1) a method that can retain the diversity of raw code to the utmost extent. 2) a LLM-based Generatior-Discriminator framework to further improve the data quality of instruction instances.
2.2.1 Raw Code Collection
To ensure the quality and diversity of raw code, we manually define some filtering rules and utilize a cluster method KCenterGreedy [28, 29] to get the raw code collection from the open source code dataset for instruction data generation. In this work, we select CodeSearchNet  [^2] https://huggingface.co/datasets/code_search_net which contains 2 million (comment, code) pairs from open-source libraries hosted on GitHub and includes code and documentation for 6 programming languages as our foundation dataset and develop a series of manually filtering rules to ensure the basic data quality of raw code. Subsequently, we apply the coreset-based selection method KCenterGreedy to maximize the diversity of the raw code. More details are as follows:
Manually defined filtering rules. In order to select high-quality code for instruction-tuning, we make some rules to filter the foundation dataset preliminarily. The main rules are as follows:
  1. i)
@@ 2.2.1 Raw Code Collection
Coreset-based selection method. After filtered by the above rules, the size of the raw code dataset dropped to 1.2 million, which is still a huge scale and too expensive to fine-tune. To this end, we employed KCenterGreedy algorithm based on code embeddings to reduce the amount of training data while maximizing data diversity of raw code as much as possible.
@@
Table 2: The proportion of generated data and the prompts used in generation phase.
@@
Figure 3: The overview of the LLM-based Generator-Discriminator Framework. In part A, the output of Generator includes 4 keys: Task name, Instruction, Information, Solution. All keys will be analyzed in the Discrimination Phase.
@@ 2.2.2 LLM-based Generator-Discriminator Framework
After the process of collecting raw code, it is more significant to generate high-quality and diverse instruction data and reformat them to training data. To this end, shown in Figure 2, we propose a novel LLM-based Generator-Discriminator Framework, where the Generator can leverage an extensive amount of unsupervised open source code to generate supervised instruction data. and the Discriminator can continue analyzing and filtering the instruction data. By incorporating such a framework into the open source code, the diversity of the generated data no longer relies solely on capability of the teacher LLM itself and the data quality of instruction data generation can be enhanced by the Discriminator.
Figure 4: An example of LLM-based Generator and Discriminator framework on code generation task. The generator produce the instruction data based on the input (a). Subsequently, the discriminator accept the output and generated analysis for it. The output (b) includes four keys, we take the information as input and solution as output in our instruction tuning. The analysis (c) consists of the detailed reason for each rule and an overall answer to check if the sample meet all the requirements.
Generation Phase. In the generation phase, we utilize GPT-4 to generate a task definition in the context of a specific scenario. As shown in Figure 3, following the model-generated task definition, we manually develop the generation requirements for the each code-related task. Integrating both the task definition and all associated requirements into the generation prompt, we take the raw code as an input and select good and bad case examples from the example database to leverage the few-shot abilities of GPT-3.5 to generate the necessary knowledge and information for instruction tuning. However, based on our experiments, we observe that few-shot generation is not always the best way to get instruction data especially on code generation task. We analyzed that this is because the teacher LLM always imitated the given example during the few-shot generation process, which will damage the diversity of instruction data. Therefore, we employed zero-shot generation on code generation task and few-shot generation on other tasks.
Discrimination Phase. During the exploration of the instruction generation process, we noticed that the data quality of the instruction instances cannot be guaranteed through the generation phase alone. In order to enhance the controllability of data generation and further ensure the data quality, we designed a LLM-based discriminator to continue analyzing and filtering the instruction data. For the discrimination phase, we establish a series of criteria related to instruction data and employ GPT-4 to assess the quality of instruction instances. However, due to the large number of rules, the discriminator cannot conduct effective and comprehensive discrimination. To solve this problem, we disassemble the discrimination rules to some subtopics inspired by Zero-shot-CoT [30] and employ the LLM-based discriminator to analyze the generation step by step . By adopting this method, the discrimination process can be more contrallable when filtering low-quality instruction instances. After the discrimination process, each instruction instance is classified as either a good or bad case and the classification information is subsequently reused in the following generation as examples. For the reusage of these classified instruction instance, different from self-instruct [31] which solely utilize the exemplification as good example, we exploit both the good generation and bad generation as the few-shot example so that the generator can gain from the analysis of bad example. An example of our Generator and Discriminator framework is presented in Figure 4. This framework provides a comprehensive approach to generating and evaluating instruction data, ensuring a high-quality and diverse training dataset.
3 Experiments
3.1 Training
@@
To obtain WaveCoder models, We choose StarCoder-15B, CodeLLaMa (7B and 13B), DeepseekCoder-6.7B as the base model and fine-tune all the base model for 3 epochs using NVIDIA A100-80GB GPU. For StarCoder-15B and CodeLLaMa-13B, we set the global batch size to 128 using Tensor Parallel and set the initial learning rate at 2e-5. For DeepseekCoder-6.7B and CodeLLaMa-7B, we set the global batch size to 256 using the Fully Sharded Data Parallel (FSDP) module from Pytorch and set the initial learning rate at 5e-5. The sequence length of our each instruction sample is less than 1024. Followed Alpaca [12], we set the fine-tuning prompt as follows:
@@
3.2 Evaluation
@@
DeepseekCoder [22] 6.7B - 49.4 60.6
Magicoder-DS[34] 6.7B DeepseekCoder 75K 66.5 60.4
WaveCoder-DS-6.7B [Uncaptioned image] 6.7B DeepseekCoder 20K 64.0 (+14.6) 62.8 (+2.2)
3.2.1 Benchmarks and Baselines
MBPP  [^5] https://huggingface.co/datasets/mbpp consists of around 1,000 crowd-sourced Python programming problems, designed to be solvable by entry level programmers, covering programming fundamentals, standard library functionality, and so on. In this paper, we choose the 500 problems test dataset to evaluate both few-shot inference of fine-tuned models. For whose MBPP (500) result is not reported or not used, we reproduced for them using bigcode-evaluation-harness  [^6] https://github.com/bigcode-project/bigcode-evaluation-harness.
Proprietary Models. We present the self-reported results from an array of SoTA LLMs, including ChatGPT (gpt-3.5-turbo), GPT-4. If not reported, we use the results from Octopack [21] or evaluate by ourselves.
3.2.2 metrics
$pass@k=E\left[1-\frac{{n\choose k-c}}{{n\choose k}}\right]$ (1)
3.2.3 Evaluation on Code Generation Task
WaveCoder models substantially outperform the instructed models training with fewer than 20K instruction tuning data (InsT Data) and CodeAlpaca with 20K data. Following the fine-tuning process, the performance of our models exhibit substantial improvement when compared to both the foundation model and a selection of open-source models, but it still lags behind proprietary models the instructed models training with more than 70K training data.
Refined and diverse instruction data can significantly improve the efficiency of instruction tuning. According to the comparison in Table 3, WaveCoder solely relies on a dataset encompassing 20,000 entries to accomplish performance levels on par with those of Magicoder [34] and Wizardcoder [14]. Although WaveCoder falls behind WizardCoder [14] and Magicoder [34] on code generation task, there is a considerable difference in data size among them. Nevertheless, the performance lag of WaveCoder is relatively minor.
@@ 3.2.3 Evaluation on Code Generation Task
WaveCoder-DS-6.7B [Uncaptioned image] 57.9 52.4 57.3 47.5 45.1 36.0 49.4
@@ 3.2.4 Evaluation on Other Code-related Task
We score WaveCoder with state-of-the-art Code LLMs on HumanEvalPack in Table 4, 5 and use the result reported in Octopack [21] when avaliable. Table 4 presents the results of WaveCoder on code repair task and Table 5 presents the results of WaveCoder on Code summarization task, highlighting the the following salient observations:
@@
Magicoder-DS 55.5 36.6 49.4 36.0 39.6 27.4 40.7
WaveCoder-DS-6.7B [Uncaptioned image] 48.2 47.5 49.4 32.3 48.2 22.0 41.3
@@ 3.3 Comparison with CodeAlpaca
1) After being fine-tuned with 1K, 5K and 20K of instructional data respectively, the performance of base model improves significantly on HumanEval shown in Figure 1. Taking Starcoder as the base model, CodeOcean surpasses the CodeAlpaca (44.9% vs 41.7%, 45.7% vs 48.1% and 47.0% vs 50.5%) shown in Figure 1 (a), which emphasizes the effectiveness of our LLM-based Generator-Discriminator Framework on refining instruction data. As shown in Figure 1 (b), The results of different base models on CodeOcean surpasses the results on CodeAlpaca, which emphasizes the effectiveness of CodeOcean dataset in enhancing the instruction-following ability of the base model.
2) According to Table 4 and Table 5, All WaveCoder models significantly outperform the model fine-tuned with CodeAlpaca. Remarkably, The pass@1 score of WaveCoder-CL-13B outperforms CodeLLaMa-CodeAlpaca-13B achieving 10.6% absolute improvements on HumanEvalExplain. This emphasizes the effectiveness of defining and classifying code-related tasks on enhancing the generalization ability of Code LLMs.
4 Ablation and Analysis
To evaluate the the effectiveness of each code-related task appearing in training data and the relationship between different tasks, we conduct an ablation study about the task type of instruction data. Using DeepseekCoder-Base-6.7B as our base model, we have the following observations from Table 6:
Refined instruction data can significantly improve the generalization ability of pre-trained models without a tradeoff. As shown in Table 6, incorporating all 4 code-related tasks into training data, WaveCoder-DS-6.7B achieves the best performance on benchmark of all tasks. For example, the participation of the Code Repair task yields a considerable average improvement of 33.7% absolute for HumanEvalFix without any significant decline in other tasks, and even improved by 3.1% absolute for HumanEval benchmark.
Different tasks can promote each other so that the model can show a generalization ability. From Table 6, we can observe that any combination of three tasks resulted in a lower score than all tasks. For example, the addition of the code summarization task offers a modest yet significant average improvement on all benchmarks. Moreover, the absence of any task will cause the score of HumanEval to drop, which also reflects the mutual promotion between different tasks.
Table 6: Ablation study on different code-related tasks: CG (Code Generation), CS (Code Summarization), CT (Code Translation), CR (Code Repair). WaveCoder-DS-6.7B utilizes all 4 code-related tasks.
Model CG CS CT CR HumanEval HumanEval Fix (Avg.) HumanEval Explain (Avg.)
DeepseekCoder-Base-6.7B 49.4 29.0 34.6
WaveCoder-DS-6.7B [Uncaptioned image] 64.0 (+14.6) 49.4 (+20.4) 41.3 (+7.3)
-Without Code Repair 60.9 (-3.1) 15.7 (-33.7) 41.2 (-0.1)
-Without Code Generation 53.6 (-10.4) 47.4 (-2.0) 40.5 (-0.8)
-Without Code Translation 60.9 (-3.1) 49.3 (-0.1) 41.6 (+0.3)
-Without Code Summarization 61.5 (-2.5) 45.6 (-3.8) 28.4 (-12.9)
@@ 5 Related Work
5.1 Code Large Language Models
Numerous researchers have proposed their Code Large Language Models (LLMs), including CodeGen [36], CodeT5[37], StarCoder [6], CodeLLaMa [7], and Deepseek-Coder [22] within coding domain. By leveraging a comprehensive pre-training process with a vast amount of code knowledge from scratch, these Code LLMs often exceed the performance of super LLMs in many scenarios. In a bid to generate output more efficiently and assist in resolving programming issues, instruction-based models like InstructCodeT5+ [38] and WizardCoder [14] have been introduced. When drafting this paper, we noticed that our contemporaneous work Magicoder [34] has been released, so we conduct a brief comparative analysis in Section 3.
5.2 Instruction Tuning
In order to unleash the potential of the foundation LLMs, instruction tuning is proposed to fine-tune the LLM with multi-task datasets which is presented in task descriptions by natural language. For example, OpenAI fine-tuned its GPT-3 [3] model with a diverse instruction dataset written by human annotators, the instruction-tuned model can effectively align the intention of human users. This work is the indispensable part of the InstructGPT [17], which is the foundation of the well-known ChatGPT. Flan-PaLM 540B [10] is instruction-tuned on 1.8K tasks and finds that the diversity of instruction dataset can significantly improve the performance of the instruction-tuned language model. Alpaca [12] adopts the self-instruct method, which leverages the few shot ability of ChatGPT to generate data for training. For Code LLMs, InstructCodeT5+ [38], WizardCoder [14], Pangu-coder2 [32], are recent recent models trained with coding instructions. However, all the instruction data they used is from CodeAlpaca, which is not refined enough.